OpenStudy (anonymous):

3(x – 4) = 33 Prove: x = 15 3x – 12 = 33 would be an example of what?

OpenStudy (johnweldon1993):

Are we solving both of those questions?

hero (hero):

Is this about properties as well?

hero (hero):

I would have divided both sides by 3 to get x - 4 = 11

OpenStudy (anonymous):

it's the same thing I just added a bit more to it

OpenStudy (johnweldon1993):

Well Prove x = 15 Just plug it in for x in the equation 3(x - 4) = 33 3(15 - 4) = 33 3(11) = 33 33 = 33 so that is proven... Now....what is THAT an example of? going from 3(x - 4) to 3x - 12 Is called the distributive property...when you apply the number on the outside of the parenthesis to each term in the parenthesis

OpenStudy (anonymous):

Thankyou :D

OpenStudy (johnweldon1993):

No problem :)